OpenStudy (liddoislander):

Find the slope of line passing through the points (4,8) and (6, -2). A. 1/5 B. 5 C. -1/5 D. -5

jimthompson5910 (jim_thompson5910):

Use the slope formula \[\Large m = \frac{y_2 - y_1}{x_2-x_1}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) = (4,8) (x2,y2) = (6,-2)

OpenStudy (liddoislander):

@jim_thompson5910 i dont get it :/

jimthompson5910 (jim_thompson5910):

\[\Large m = \frac{y_2 - y_1}{x_2-x_1}\] \[\Large m = \frac{-2-8}{6 - 4}\] \[\Large m = ???\]

OpenStudy (liddoislander):

-1/5?

OpenStudy (liddoislander):

1/5

jimthompson5910 (jim_thompson5910):

-2-8 is equal to?